References

References

R. Abraham and J. Marsden, Foundations of Mechanics, Benjamin/Cummings, Reading, Mas-

sachusetts, 1978. J. Adams, Lectures on Lie Groups, W. A. Benjamin, New York, 1969. J. Adams, D. Barbasch, and D. Vogan, The Langlands Classification and Irreducible Characters

for Real Reductive Groups, Birkhäuser, Boston-Basel-Berlin, 1992. J. Adams, J.-S. Huang, and D. Vogan, Functions on the model orbit in E_{8}, Represent. Theory 2

(1998), 224–263. J. Adams and D. Vogan, Harish-Chandra’s method of descent to appear. J. Adams and D. Vogan, L-groups, projective representations, and the Langlands classification,

Amer. J. Math. 114 (1992), 45–138. J. Adams and D. Vogan, Lifting of characters and Harish-Chandra’s method of descent, to appear. M. Andler, Relationships of divisibility between local L-functions associated to representations of

complex reductive groups.”. A. Andreotti and H. Grauert, Théorèmes de finitude pour la cohomologie des espaces complexes,

Bull. Soc. Math. France 90 (1962), 193–259. V. Arnold, Mathematical Methods of Classical Mechanics, Springer-Verlag, New York-Heidelberg-

Berlin, 1978. J. Arthur, Eisenstein series and the trace formula”, Automorphic Forms, Representations, and

L-functions, Proceedings of Symposia in Pure Mathematics, vol. 33, part 2, American Mathematical Society, Providence, Rhode Island, 1979, pp. 27–61.

J. Arthur, On some problems suggested by the trace formula, Lie Group Representations II R. Herb, R. Lipsman, and J. Rosenberg, eds. Lecture Notes in Mathematics, vol. 1041, Springer-Verlag, Berlin-Heidelberg-New York, 1983, pp. 1–49.

J. Arthur, Unipotent automorphic representations: conjectures 13–71, Orbites Unipotentes et

Représentations II. Groupes p-adiques et Réels, Astérisque, vol. 171–172, 1989. M. Artin, Néron models, Arithmetic Geometry (G. Cornell and J. Silverman, eds.), Springer-

Verlag, New York-Berlin-Heidelberg, 1986, pp. 213–230. M. Atiyah and I. MacDonald, Introduction to Commutative Algebra Addison-Wesley, Reading,

Massachusetts, 1969. M. Atiyah and W. Schmid, A geometric construction of the discrete series for semisimple Lie

groups, Invent. Math. 42 (1977), 1–62. L. Auslander and B. Kostant, Polarization and unitary representations of solvable Lie groups,

Invent. Math. 14 (1971), 255–354. M. W. Baldoni-Silva and D. Barbasch, The unitary spectrum for real rank one groups, Invent.

Math. 72 (1983), 27–55. M. W. Baldoni-Silva and A. Knapp, Indefinite intertwining operators, Proc. Nat. Acad. Sci. U.S.A.

81 (1984), 1272–1275. M. W. Baldoni-Silva and A. Knapp, Unitary representations induced from maximal parabolic

subgroups, J. Funct. Anal. 69 (1986), 21–120. M. W. Baldoni-Silva and A. Knapp, Intertwining operators and unitary representations II, Orbites

Unipotentes et Représentations II. Groupes p-adiques et réels, Astérisque, vol. 173-174, Société Mathématique de France, Paris, 1989, pp. 131–162.

D. Barbasch, The unitary dual for complex classical Lie groups Invent. Math. 96 (1989), 103–176. D. Barbasch and A. Moy, A unitarity criterion for p-adic groups Invent. Math. 98 (1989), 19–37. D. Barbasch and A. Moy, Reduction to real infinitesimal character in affine Hecke algebras, J.

Amer. Math. Soc. 6 (1993,), 611–635. D. Barbasch and A. Moy, Unitary spherical spectrum for p-adic classical groups, Acta Appl.

Math. 44 (1996), 3–37. D. Barbasch and D. Vogan, The local structure of characters, J. Funct. Anal. 37 (1980), 27–55. D. Barbasch and D. Vogan, Weyl group representations and nilpotent orbits, Representation

Theory of Reductive Groups (P. Trombi, eds.), Birkhäuser, Boston-Basel-Stuttgart, 1983, pp. 21-33.

D. Barbasch and D. Vogan, Problems in primitive ideal theory to appear in Proceedings of the Durham Conference on Enveloping Algebras.

D. Barbasch and D. Vogan, Unipotent representations of complex semisimple Lie groups, Ann. of

Math. 121 (1985), 41–110. V. Bargmann, Irreducible unitary representations of the Lorentz group, Ann. of Math. 48 (1947),

568–640. F. Bayen, M. Flato, C. Fronsdal, A. Lichnerowicz, and D. Sternheimer, Deformation theory and

quantization. I. Deformation of symplectic structures, Ann. Physics 111 (1978), 61–110. A. Beilinson and J. Bernstein, Localisation de 𝔤-modules, C. R. Acad. Sci. Paris 292 (1981),

15–18. A. Beilinson, J. Bernstein, and P. Deligne, Faisceaux pervers 5–171, Analyse et Topologie sur les

Espaces Singuliers, volume 1 Astérisque, vol. 100, 1982. A. Beilinson, Localization of representations of reductive Lie algebras 699–710, Proceedings of

the International Congress of Mathematicians, Warsaw 1983 Amsterdam-London, 1984. F. A. Berezin, Laplace operators on semi-simple Lie groups, Trudy Moskov. Math. Obšč 6 (1957),

371–463; English translation:, Amer. Math. Soc. Transl. (2) 21 (1962), 239–339. N. Berline and M. Vergne, The equivariant index and Kirillov’s character formula Amer. J. Math.

107 (1985), 1159–1190. P. Bernat and J. Dixmier, Sur le dual d’un groupe de Lie C. R. Acad. Sci. Paris 250 (1960),

1778–1779. J. Bernstein, P. Deligne, and D. Kazhdan, Trace Paley-Wiener theorem for reductive p-adic groups

J. Analyse Math. 47 (1986), 180–192. J. Bernstein, rédigé par P. Deligne, Le ‘centre’ de Bernstein, Représentations des Groupes Re-

ductifs sur un Corps Local (J. Bernstein, P. Deligne, D. Kazhdan, and M.-F. Vigneras, eds.), Hermann, Paris, 1984.

J. Bernstein, P. Deligne, D. Kazhdan, and M.-F. Vigneras, Représentations des Groupes Reductifs

sur un Corps Local, Hermann, Paris, 1984. I. N. Bernstein, I. M. Gelfand, and S. I. Gelfand, Structure of representatations generated by

vectors of highest weight, Funct. Anal. Appl. 5 (1972), 1–8. I. N. Bernstein, I. M. Gelfand, and S. I. Gelfand, Models of representations of compact Lie groups,

Funct. Anal. Appl. 9 (1975), 61–62. J. Bernstein and A. V. Zelevinsky, Induced representations of reductive p-adic groups I Ann. Sci.

Ècole Norm. Sup. 10 (1977), 441–472. F. Bien, Spherical D-modules and Representations of Reductive Lie Groups, Ph.D. dissertation,

Massachusetts Institute of Technology, Cambridge, Massachusetts, 1986. H. Boerner, Representations of Groups. With special consideration for the needs of modern

physics, American Elsevier Publishing Co., Inc., New York, 1970. A. Borel, Admissible representations of a semisimple group over a local field with fixed vectors

under an Iwahori subgroup, Invent. Math. 35 (1976), 233–259. A. Borel, Automorphic L-functions, Automorphic Forms, Representations, and L-functions, Pro-

ceedings of Symposia in Pure Mathematics, vol. 33, part 2, American Mathematical Society, Providence, Rhode Island, 1979, pp. 27–61.

A. Borel, Linear Algebraic Groups, Springer-Verlag, New York, 1991. A. Borel et al., Intersection Cohomology, Birkhäuser, Boston-Basel-Stuttgart, 1984. A. Borel et al., Algebraic D-Modules, Perspectives in Mathematics, vol. 2, Academic Press, Boston,

1987. A. Borel and J.-P. Serre, Détermination des p-puissances réduites de Steenrod dans la cohomologie

des groupes classiques. Applications, C. R. Acad. Sci. Paris 233 (1951), 680–682. A. Borel and N. Wallach, Continuous Cohomology, Discrete Subgroups, and Representations of

Reductive Groups, Princeton University Press, Princeton, New Jersey, 1980. W. Borho, Primitive vollprime Ideale in der Einhüllenden von 𝔰𝔬(5, ℂ), J. Algebra 43 (1976),

619–654. W. Borho, Definition einer Dixmier-Abbildung für 𝔰l(n, ℂ), Invent. Math. 40 (1977), 143–169. W. Borho, Über Schichten halbeinfacher Lie-Algebren Invent. Math. 65 (1981), 283-317. W. Borho, A survey on enveloping algebras of semisimple Lie algebras, Lie Algebras and Related

Topics (D. Britten, F. Lemire, and R. Moody, eds.), CMS Conference Proceedings, volume 5, American Mathematical Society for CMS, Providence, Rhode Island, 1986.

W. Borho and J.-L. Brylinski, Differential operators on homogeneous spaces I, Invent. Math. 69 (1982), 437–476.

W. Borho and J.-L. Brylinski, Differential operators on homogeneous spaces III, Invent. Math.

80 (1985), 1–68. W. Borho, P. Gabriel, and R. Rentschler, Primideale in Einhüllenden auflosbarer Lie-Algebren,

Lecture Notes in Mathematics, vol. 357, Springer-Verlag, Berlin-Heidelberg-New York, 1973. W. Borho and J. C. Jantzen, Über primitive Ideale in der Einhüllenden einer halbeinfachen Lie-

Algebra, Invent. Math. 39 (1977), 1–53. W. Borho and H. Kraft, Über Bahnen und deren Deformationen bei linearen Aktionen reduktiver

Gruppen, Comment. Math. Helvetici 54 (1979), 61–104. R. Bott, Homogeneous vector bundles, Ann. of Math. 66 (1957), 203–248. N. Bourbaki, Groupes et Algèbres de Lie. Chapitres 4, 5, et 6, Masson, Paris, 1981. N. Bourbaki, Topological vector spaces. Chapters 1–5, Springer-Verlag, Berlin-Heidelberg-New

York, 1987. T. Bratten, Realizing representations on generalized flag manifolds, Compositio math. 106 (1997),

283–319. T. Bratten, A comparison theorem for Lie algebra homology groups, Pacific J. Math. 182 (1998),

23–36. T. Bröcker and T. tom Dieck, Representations of compact Lie groups, Springer-Verlag, Berlin-

Heidelberg-New York, 1985. R. Brylinski and B. Kostant, Nilpotent orbits, normality, and Hamiltonian group actions, preprint. F. Bruhat, Sur les représentations induites des groupes de Lie, Bull. Soc. Math. France 84 (1956),

97–205. J.- L. Brylinski and M. Kashiwara, Kazhdan-Lusztig conjecture and holonomic systems, Invent.

Math. 64 (1981), 387–410. M. Burger, J. S. Li, and P. Sarnak, Ramanujan duals and the automorphic spectrum, Bull. Amer.

Math. Soc. (N.S.) 26 (1992), 253–257. C. Bushnell and P.Kutzko, Smooth representations of reductive p-adic groups: structure theory

via types, Proc. London Math. Soc. (3) 77 (1998), 582–634. J. Carmona, Sur la classification des modules admissibles irréductibles 11–34, Non-commutative

Harmonic Analysis and Lie Groups (J. Carmona and M. Vergne, eds.), Lecture Notes in Mathematics, vol. 1020, Springer-Verlag, Berlin-Heidelberg-New York-Tokyo, 1983.

É. Cartan, La Théorie des Groupes Finis et Continus et l’Analysis Situs, Gauthier-Villars, Paris,

1930. É. Cartan, Sur la Structure des Groupes de Transformations Finis et Continus, Libraire Vuibert,

Paris, 1933. H. Cartan and S. Eilenberg, Homological Algebra, Princeton University Press, Princeton, New

Jersey, 1956. R. Carter, Finite Groups of Lie type. Conjugacy Classes and Complex Characters, John Wiley &

Sons Ltd, Chichester, England, 1985. W. Casselman, Canonical extensions of Harish-Chandra modules to representations of G Can. J.

Math. 41 (1989), 385–438. J. Cassels and A. Fröhlich, Algebraic Number Theory, Academic Press, London, 1967. P. Chernoff, Mathematical obstructions to quantization. Proceedings of the Third Workshop on

Lie-Admissible Formulations (Univ. Massachusetts, Boston, Mass., 1980), Part B, Hadronic J. 4 (1980/81 pages 879–898).

C. Chevalley, Theory of Lie Groups, Princeton University Press, Princeton, New Jersey, 1946. C. Chevalley, The Algebraic Theory of Spinors, Columbia University Press, New York, 1954. N. Chriss and V. Ginzburg, Representation Theory and Complex Geometry, Birkhäuser, Boston-

Basel-Berlin, 1997. D. Collingwood and W. McGovern, Nilpotent Orbits in Semisimple Lie Algebras, Van Nostrand

Reinhold, New York, 1993. N. Conze, Quotients primitifs des algèbres enveloppantes et algèbres d’opérateurs différentiels, C.

R. Acad. Sci. Paris Sér A-B 277 (1973), A1033-A1036. N. Conze and J. Dixmier, Ideaux primitifs dans l’algèbre enveloppante d’une algèbre de Lie semi-

simple, Bull. Sci. Math.(2) 96 (1972), 339–351. N. Conze-Berline and M. Duflo, Sur les représentations induites des groupes semi-simples com-

plexes, Compositio math. 34 (1977), 307–336. C. Curtis, N. Iwahori, and R. Kilmoyer, Hecke algebras and characters of parabolic type of finite

groups with (B, N)-pairs, Inst. Hautes Études Sci. Publ. Math. 40 (1971), 81–116.

P. Deligne and G. Lusztig, Representations of reductive groups over finite fields, Ann. of Math.

103 (1976), 103–161. M. Demazure, A. Grothendieck, it et al., Schémas en Groupes I (SGA 3), Lecture Notes in

Mathematics, vol. 151, Springer-Verlag, Berlin-Heidelberg-New York, 1970. M. Demazure, A. Grothendieck, it et al., Schémas en Groupes III (SGA 3), Lecture Notes in

Mathematics, vol. 153, Springer-Verlag, Berlin-Heidelberg-New York, 1970. M. De Wilde and P. Lecomte, Existence of star-products and formal deformations of the Poisson

Lie algebra of arbitrary symplectic manifolds, Lett. Math. Phys. 7 (1983), 487–496. J. Dixmier, Les C^{∗}-algèbres et leurs Représentations, Gauthier-Villars, Paris, 1964. J. Dixmier, Von Neumann Algebras, translated by F. Jellett, North-Holland Publishing Co., New

York, 1981. J. Dixmier, Algèbres Enveloppantes, Gauthier-Villars, Paris-Brussels-Montreal, 1974. J. Dixmier, Polarisations dans les algèbres de Lie semi-simples complexes, Bull. Sci. Math.(2) 99

(1975), 45–63. J. Dixmier, Ideaux primitifs dans les algèbres enveloppantes, J. Algebra 48 (1978), 96–112. J. Dixmier and P. Malliavin, Factorisations de fonctions et de vecteurs indéfiniment différentiables,

Bull. Sci. Math. (2) 102 (1978),), 307–330. D. Doković, Classification of nilpotent elements in simple exceptional real Lie algebras of inner

type and description of their centralizers, J. Algebra 112 (1988), 503–524. D. Doković, Classification of nilpotent elements in simple real Lie algebras E_{6(6)} and E_{6(−26)}

and description of their centralizers, J. Algebra 116 (1988), 196–207. M. Duflo, Représentations irréductibles des groupes semi-simples complexes, Analyse Harmonique

sur les Groupes de Lie (Sém. Nancy-Strasbourg 1973–75) Lecture Notes in Mathematics, vol. 497, Springer-Verlag, Berlin-Heidelberg-New York-Tokyo, 1975, pp. 26–88.

M. Duflo, Sur la classification des idéaux primitifs dans l’algèbre enveloppante d’une algèbre de

Lie semi-simple, Ann. of Math. 105 (1977), 107–120. M. Duflo, Représentations unitaires irréductibles des groupes semi-simples complexes de rang

deux, Bull. Soc. Math. France 107 (1979), 55–96. M. Duflo, Représentations unitaires des groupes semi-simples complexes, Group Theoretical Meth-

ods in Physics, Proc. Eighth Internat. Colloq., Kiryat Anavim 1979, Ann. Israel Phys. Soc., vol. 3, Hilger, Bristol, 1980, pp. 19–34.

M. Duflo, Théorie de Mackey pour les groupes de Lie algébriques, Acta Math. 149 (1982), 153–

213. M. Duflo, Sur les idéaux induits dans les algèbres enveloppantes Invent. Math. 67 (1982), 385-393. M. Duflo, Construction de représentations d’un groupe de Lie, Cours d’été du C.I.M.E., Cortona

(1980). M. Duflo, Construction de gros ensembles de représentations unitaires irréductible d’un groupe de

Lie quelconque, Operator Algebras and Group Representations, volume I, Monographs Stud. Math., vol. 17, Pitman, Boston, Massachusetts-London, 1984.

M. Duflo, G. Heckman, and M. Vergne, Projection d’orbites, formule de Kirillov et formule de Blattner, Harmonic analysis on Lie groups and symmetric spaces (Kleebach, 1983), Mém. Soc. Math. France (N.S.) No. 15, 1984, pp. 65–128.

E. B. Dynkin, Semisimple subalgebras of semisimple Lie algebras, Amer. Math. Soc. Transl., Ser.

2 6 (1957), 111–245. D. Eisenbud, Commutative algebra with a view toward algebraic geometry, Springer-Verlag, New

York Berlin Heidelberg, 1995. T. J. Enright, Relative Lie algebra cohomology and unitary representations of complex Lie groups,

Duke Math. J. 46 (1979), 513–525. T. J. Enright and N. R. Wallach, Notes on homological algebra and representations of Lie algebras,

Duke Math. J. 47 (1980), 1–15. M. Flensted-Jensen, Discrete series for semisimple symmetric spaces, Ann. of Math. 111 (1980),

253–311. O. Gabber, The integrability of the characteristic variety, Amer. J. Math. 103 (1981), 445–468. D. Garfinkle, The annihilators of irreducible Harish-Chandra modules for SU(p, q) and other type

A_{n−1} groups, to appear. E. Garnica Vigil, On the decomposition of Langlands subrepresentations for a group in the Harish-

Chandra class, Trans. Amer. Math. Soc. 347 (1995), 1609–1648.

E. Garnica Vigil, An action of the R-group on the Langlands subrepresentations, Geometry and Representation Theory of Real and p-adic Groups (J. Tirao, D. Vogan, and J. Wolf, eds.), Birkhäuser, Boston-Basel-Berlin, 1998, pp. 57–67.

I. M. Gelfand, Some aspects of functional analysis and algebra, Proceedings of the International Congress of Mathematicians 1954 (Amsterdam), vol. I, Erven P. Noordhoof, N. V., Groningen, North Holland Publishing Co., Amsterdam, 1957, pp. 253–276.

I. M. Gelfand and M. A. Naimark, Unitary representations of the Lorentz group, Izv. Akad. Nauk

S.S.S.R. 11 (1947), 411–504. I. M. Gelfand and M. A. Naimark, Unitary Representations of the Classical Groups, Trudy Mat.

Inst. Steklov, vol. 36 Moscow-Leningrad, 1950; German translation:, Akademie-Verlag, Berlin, 1957.

S. I. Gelfand, Weil’s representation of the Lie algebra of type G_{2}, and representations of SL_{3}

connected with it, Funct. Anal. Appl. 14 (1980), 40–41. M. Gerstenhaber, On the deformation of rings and algebras, Ann. of Math. 79 (1964), 59–103. V. Ginsburg, Characteristic varieties and vanishing cycles, Invent. Math. 84 (1986), 327–402. V. Ginsburg, 𝔊-modules, Springer’s representations and bivariant Chern classes, Adv. in Math.

61 (1986), 1–48. R. Godement, Topologie Algébrique et Théorie des Faisceaux, Hermann, Paris, 1958. M. Goresky and R. MacPherson, Stratified Morse Theory, Springer-Verlag, Berlin-Heidelberg-New

York, 1983. M. Goresky and R. MacPherson, Local contribution to the Lefschetz fixed point formula, preprint. W. Graham, Functions on the universal cover of the principal nilpotent orbit, Invent. Math. 108

(1992), 15–27. W. Graham and D. Vogan, Geometric quantization for nilpotent coadjoint orbits, Geometry and

Representation Theory of Real and p-adic Groups (J. Tirao, D. Vogan, and J. Wolf, eds.), Birkhäuser, Boston-Basel-Berlin, 1998, pp. 69–137.

W. Greub, S. Halperin, and R. Vanstone, Connections, Curvature, and Cohomology. Volume III: Cohomology of Principal Bundles and Homogeneous Spaces, Pure and Applied Mathematics, vol. 47-III, Academic Press (Harcourt Brace Jovanovich, Publishers), New York-London, 1976. :

H. J. Groenewold, On the principles of elementary quantum mechanics, Physica 12 (1946), 405–

460. A. Grothendieck (rédigés avec la collaboration de J. Dieudonné), Eléments de Géométrie Algébrique

IV. Etude locale des schémas et des morphismes de schémas (Seconde Partie), Publications Mathématiques, vol. 24, Institut des Hautes Etudes Scientifiques, Le Bois-Marie, Bures-sur-Yvette, France, 1965.

A. Guichardet, Théorie de Mackey et méthode des orbites selon M. Duflo, Expo. Math. 3 (1985),

303–346. A. Guichardet, Cohomologie des Groupes Topologiques et des Algèbres de Lie, CEDIC, Paris,

1980. V. Guillemin and S. Sternberg, A generalization of the notion of polarization Ann. Glob. Analysis

and Geometry 4 (1986), 327–347. V. Guillemin and S. Sternberg, Geometric Asymptotics, revised edition, Mathematical Surveys

and Monographs, vol. 14, American Mathematical Society, Providence, Rhode Island, 1990. A. Guillemonat, Représentations spheriques singulieres, Non-commutative Harmonic Analysis

and Lie Groups (J. Carmona and M. Vergne, eds.), Lecture Notes in Mathematics, vol. 1020, Springer-Verlag, Berlin-Heidelberg-New York-Tokyo, 1983.

Harish-Chandra, Representations of a semisimple Lie group on a Banach space I, Trans. Amer.

Math. Soc. 75 (1953), 185–243. Harish-Chandra, The Plancherel formula for complex semisimple Lie groups Trans. Amer. Math.

Soc. 76 (1954), 485–528. Harish-Chandra, Representations of semisimple Lie groups, V, Amer. J. Math. 78 (1956), 1–41. Harish-Chandra, Discrete series for semi-simple Lie groups I. Construction of invariant eigendis-

tributions, Acta Math. 113 (1965), 241–318. Harish-Chandra, Discrete series for semi-simple Lie groups II. Explicit determination of the

characters, Acta Math. 116 (1966), 1–111. Harish-Chandra, Harmonic analysis on real reductive groups I. The theory of the constant term,

J. Func. Anal. 19 (1975), 104–204.

R. Hartshorne, Algebraic Geometry, Springer-Verlag, New York, Heidelberg, Berlin, 1977. M. Hashizume, Whittaker models for real reductive groups Japan. J. Math. 5 (1979), 349–401. H. Hecht, D. Miličić, W. Schmid, and J. Wolf, Localization and standard modules for real semisim-

ple Lie groups I: The duality theorem, Invent. Math. 90 (1987), 297–332. G. Heckman, Projections of orbits and asymptotic behaviour of multiplicities for compact con-

nected Lie groups, Invent. Math. 67 (1982), 333–356. S. Helgason, Differential Geometry, Lie Groups, and Symmetric Spaces, Academic Press, New

York, San Francisco, London, 1978. S. Helgason, Groups and Geometric Analysis, Academic Press, Orlando, Florida, 1984. S. Helgason, Geometric Analysis on Symmetric Spaces, Mathematical Surveys and Monographs,

vol. 39, American Mathematical Society, Providence, Rhode Island, 1994. S. Helgason and K. Johnson, The bounded spherical functions on symmetric spaces, Advances in

Math. 3 (1969), 586–593. R. Herb, Fourier inversion and the Plancherel theorem for semisimple real Lie groups, Amer.

J. Math. 104 (1982), 9–58. R. Hermann, Toda lattices, cosymplectic manifolds, Bä cklund transformations and kinks. Part

A., Interdisciplinary Mathematics, Vol. XV, Math. Sci. Press, Brookline, Mass., 1977. W. Hesselink, Polarizations in the classical groups, Math. Z. 160 (1978), 217–234. V. Hinich, On the singularities of nilpotent orbits, Israel J. Math 73 (1991), 297–308. L. van Hove, Sur certaines représentations unitaires d’un groupe infini de transformations, Acad.

Roy. Belgique. Cl. Sci. Mém. Coll. in 8^{∘} 26, fasc 6 (1951), 1–102. L. van Hove, Sur le probléme des relations entre les transformations unitaires de la mécanique

quantique et les transformations canoniques de la mécanique classique, Acad. Roy. Belgique. Bull. Cl. Sci. (5) 37 (1951), 610–620.

R. Howe, On the role of the Heisenberg group in harmonic analysis, Bull. Amer. Math. Soc. (N.S.)

3 (1980), 821–843. R. Howe, Wave front sets of representations of Lie groups, Automorphic Forms, Representation

Theory, and Arithmetic, Tata Inst. Fund. Res. Studies in Math., vol. 10, Tata Institute for Fundamental Research, Bombay, 1981.

R. Howe, The oscillator semigroup, The Mathematical Heritage of Hermann Weyl (R. O. Wells, Jr., eds.), Proceedings of Symposia in Pure Mathematics, vol. 48, American Mathematical Society, Providence, Rhode Island, 1988, pp. 61–132.

R. Howe, Transcending classical invariant theory, J. Amer. Math. Soc. 2, 535–552. J.-S. Huang, The unitary dual of the universal covering group of GL(n, ℝ), Duke Math. J. 61

(1990), 705–745. J. E. Humphreys, Introduction to Lie Algebras and Representation Theory, Springer-Verlag,

Berlin-Heidelberg-New York, 1972. J. C. Jantzen, Zur Charakterformel gewisser Darstellungen halbeinfacher Gruppen und Lie-

algebren, Math. Z. 140 (1974), 127–149. J. C. Jantzen, Moduln mit einem Höchsten Gewicht, Lecture Notes in Mathematics, vol. 750,

Springer-Verlag, Berlin-Heidelberg-New York, 1979. J. C. Jantzen, Einhüllende Algebren halbeinfacher Lie-algebren, Springer-Verlag, Berlin-Heidelberg-

New York, 1983. A. Joseph, The minimal orbit in a simple Lie algebra and its associated maximal ideal, Ann. Sci.

Ecole Norm. Sup. (4) 9 (1976), 1–30. A. Joseph, Dixmier’s problem for Verma and principal series modules J. London Math. Soc.(2)

20 (1979), 193-204. A. Joseph, Goldie rank in the enveloping algebra of a semisimple Lie algebra I, II, J. Alg. 65

(1980), 269–316. A. Joseph, W-module structure in the primitive spectrum of a semisimple Lie algebra, Non-

commutative Harmonic Analysis and Lie Groups (J. Carmona and M. Vergne, eds.), Lecture Notes in Mathematics, vol. 728, Springer-Verlag, Berlin-Heidelberg-New York, 1979, pp. 116– 135.

A. Joseph, Kostant’s problem and Goldie rank, Non-commutative Harmonic Analysis and Lie Groups (J. Carmona and M. Vergne, eds.), Lecture Notes in Mathematics, vol. 880, Springer-Verlag, Berlin-Heidelberg-New York, 1981, pp. 249–266.

A. Joseph, On the classification of primitive ideals in the enveloping algebra of a semisimple Lie algebra, Lie Group Representations I (R. Herb, R. Lipsman and J. Rosenberg Lecture Notes in Mathematics, eds.), vol. 1024, Springer-Verlag, Berlin-Heidelberg-New York, 1983, pp. 30–76.

A. Joseph, On the associated variety of a primitive ideal J. Algebra 93 (1985), 509–523. A. Joseph, The primitive spectrum of an enveloping algebra, Orbites Unipotentes et Représentations

III. Orbites et Faisceaux pervers, Astérisque, vol. 173-174, 1989, pp. 13–53. M. Kashiwara, Index theorem for a maximally overdetermined system of linear differential equa-

tions Proc. Japan Acad. 49 (1973), 803–804. M. Kashiwara, Systems of Microdifferential Equations, Progress in Mathematics, vol. 34, Birkhäuser,

Boston-Basel-Stuttgart, 1983. M. Kashiwara and T. Kawai, On holonomic systems of microdifferential equations. III –Systems

with regular singularities– Publ. RIMS, Kyoto Univ. 17 (1981), 813–979. M. Kashiwara and T. Kawai, Microlocal analysis, preprint, Research Institute for Mathematical

Sciences (1983). M. Kashiwara and P. Schapira, Microlocal Study of Sheaves, Astérisque, vol. 128, 1985. D. Kazhdan, Connection of the dual space of a group with the structure of its closed subgroups

Funct. Anal. Appl. 1 (1967), 63–65. D. Kazhdan and G. Lusztig, Representations of Coxeter groups and Hecke algebras, Invent. Math.

53 (1979), 165–184. D. Kazhdan and G. Lusztig, Schubert varieties and Poincaré duality, Geometry of the Laplace

Operator, Proceedings of Symposia in Pure Mathematics, vol. 36, American Mathematical Society, Providence, Rhode Island, 1980, pp. 185–203.

D. Kazhdan and G. Lusztig, Proof of the Deligne-Langlands conjecture for Hecke algebras, Invent.

Math. 87 (1987), 153–215. J. Kim, Hecke algebras of classical groups over p-adic fields. II, Compositio Math. 127 (2001),

117–167. R. C. King, Branching rules for classical Lie groups using tensor and spinor methods, J. Phys. A

8 (1975), 429–449. A. Kirillov, Unitary representations of nilpotent Lie groups, Uspehi Mat. Nauk. 17 (1962), 57–110. A. Kirillov, Elementy Teorii Predstavlenii, Nauka, Moscow, 1972. A. Kirillov, Elements of the Theory of Representations, translated by E. Hewitt, Springer-Verlag,

Berlin-Heidelberg-New York, 1976. A. Knapp, Weyl group of a cuspidal parabolic, Ann. Sci. Ecole Norm. Sup. 8 (1975), 275–294. A. Knapp, Representation Theory of Semisimple Groups: An Overview Based on Examples,

Princeton University Press, Princeton, New Jersey, 1986. A. Knapp, Lie Groups, Lie Algebras, and Cohomology, Mathematical Notes 34, Princeton Uni-

versity Press, Princeton, New Jersey, 1988. A. Knapp, Lie Groups Beyond an Introduction, Progress in Mathematics 140, Birkhäuser, Boston-

Basel-Berlin, 1996. A. Knapp and B. Speh, Irreducible unitary representations of SU(2, 2), J. Funct. Anal 45 (1982),

41–73. A. Knapp and B. Speh, The role of basic cases in classification: theorems about unitary rep-

resentations applicable to SU(N,2), Non-commutative Harmonic Analysis and Lie Groups (J. Carmona and M. Vergne, eds.), Lecture Notes in Mathematics, vol. 1020, Springer-Verlag, Berlin-Heidelberg-New York-Tokyo, 1983, pp. 119–160.

A. Knapp and E. Stein, Singular integrals and the principal series. IV, Proc. Nat. Acad. Sci.

U.S.A. 72 (1975), 2459–2461. A. Knapp and E. Stein, Intertwining operators for semisimple groups II, Invent. Math. 60 (1980),

9–84. A. Knapp and D. Vogan, Cohomological Induction and Unitary Representations, Princeton Uni-

versity Press, Princeton, New Jersey, 1995. A. Knapp and G. Zuckerman, Classification theorems for representations of semisimple Lie

groups, Non-commutative Harmonic Analysis (J. Carmona and M. Vergne, eds.), Lecture Notes in Mathematics, vol. 587, Springer-Verlag, Berlin-Heidelberg-New York, 1977, pp. 138– 159.

A. Knapp and G. Zuckerman, Classification of irreducible tempered representations of semisimple Lie groups Ann. of Math. 116 (1982), 389–501, See also; Ann. of Math. 119 (1984), 639.

M. Kneser, Galois-Kohomologie halbeinfacher algebraischer Gruppen über 𝔭-adischen Körpern.

II, Math. Z. 89 (1965), 250–272. K. Koike, On the decomposition of tensor products of the representations of the classical groups:

by means of the universal characters, Adv. in Math. 74 (1989), 57–86. B. Kostant, The principal three-dimensional subgroup and the Betti numbers of a complex simple

Lie group, Amer. J. Math. 81 (1959), 973–1032. B. Kostant, Lie algebra cohomology and the generalized Borel-Weil theorem, Ann. of Math. 74

(1961), 329–387. B. Kostant, Orbits, symplectic structures, and representation theory, Proceedings of the United

States-Japan Seminar in Differential Geometry, Kyoto, Japan, 1965, Nippon Hyoronsha, Tokyo, 1966, p. 71.

B. Kostant, On the existence and irreducibility of certain series of representations, Bull. Amer.

Math. Soc. 75 (1969), 627–642. B. Kostant, Quantization and unitary representations, Lectures in Modern Analysis and Appli-

cations (C. Taam, eds.), Lecture Notes in Mathematics, vol. 170, Springer-Verlag, Berlin-Heidelberg-New York, 1970, pp. 87–208.

B. Kostant, On the tensor product of a finite and an infinite dimensional representation, J. Func.

Anal. 20 (1975), 257–285. B. Kostant, On the existence and irreducibility of certain series of representations, Lie groups

and their representations (Proc. Summer School, Bolyai János Math. Soc., Budapest, 1971), Halsted, New York, 1975, pp. 231–329.

B. Kostant, Symplectic spinors, Geometria Simplettica e Fisica Matematica. Symposia Mathematica XIV, Istituto Nazionale di Alta Matematica, Academic Press, London and New York, 1974, pp. 139–152.

B. Kostant, On Whittaker vectors and representation theory, Invent. Math. 48 (1979), 101–184. B. Kostant, Coadjoint orbits and a new symbol calculus for line bundles, Conference on Differential

Geometric Methods in Theoretical Physics (G. Denardo and H. D. Doebner, eds.), World Scientific, Singapore, 1983.

B. Kostant, The principle of triality and a distinguished unitary representation of SO(4, 4), Differential Geometrical Methods in Theoretical Physics (K. Bleuler and M. Werner, eds.), Kluwer Academic Publishers, Dordrecht-Boston, 1988, pp. 65–109.

B. Kostant and S. Rallis, Orbits and representations associated with symmetric spaces, Amer.

J. Math. 93 (1971), 753–809. R. Kottwitz, Sign changes in harmonic analysis on reductive groups Trans. Amer. Math. Soc.

278 (1983), 289–297. R. Kottwitz, Stable trace formula: cuspidal tempered terms Duke Math. J. 51 (1984), 611–650. R. Kottwitz, Shimura varieties and λ-adic representations 161–209, Automorphic Forms, Shimura

Varieties, and L-functions, vol. I (L. Clozel and J. Milne, eds.), Perspectives in Mathematics, vol. 10 Academic Press, San Diego, 1990.

H. Kraft and C. Procesi, Closures of conjugacy classes of matrices are normal, Invent. Math. 53

(1979), 227–247. H. Kraft and C. Procesi, On the geometry of conjugacy classes in classical groups, Comm. Math.

Helv. 57(4) (1982), 539–602. S. Kumaresan, On the canonical 𝔨-types in the irreducible unitary 𝔤-modules with non-zero relative

cohomology, Invent. Math. 59 (1980), 1–11. J.-P. Labesse and R. P. Langlands, L-indistinguishability for SL(2) Canad. J. Math. 31 (1979),

726–785. S. Lang, SL(2, ℝ), Addison-Wesley. S. Lang, Rapport sur la Cohomologie des Groupes, W. A. Benjamin, New York, 1966. R. P. Langlands, Representations of abelian algebraic groups, preprint, Yale University (1968). R. P. Langlands, Problems in the theory of automorphic forms, Lectures in modern analysis

and applications, III, Lecture Notes in Mathematics, vol. 170, Springer-Verlag, Berlin, 1970, pp. 18–61.

R. P. Langlands, On the functional equations satisfied by Eisenstein series, Lecture Notes in

Mathematics, vol. 544, Springer-Verlag, Berlin-New York, 1976. R. P. Langlands, On the classification of representations of real algebraic groups, Representation

Theory and Harmonic Analysis on Semisimple Lie Groups (P. Sally and D. Vogan, eds.), Mathematical Surveys and Monographs, vol. 31, American Mathematical Society, Providence, Rhode Island, 1989, pp. 101–170.

R. P. Langlands, Stable conjugacy: definitions and lemmas Canad. J. Math. 31 (1979), 700–725. R. P. Langlands and D. Shelstad, On the definition of transfer factors, Math. Ann. 278 (1987),

219–271. J. Lepowsky, Algebraic results on representations of semisimple Lie groups, Trans. Amer. Math.

Soc. 176 (1973), 1–44. T. Levasseur and S. P. Smith, Primitive ideals and nilpotent orbits in type G_{2}, J. Algebra 114

(1988), 81–105. J.-S. Li, Singular unitary representations of classical groups Invent. Math. 97 (1989), 237–255. Jian-shu Li, On the first eigenvalue of Laplacian for locally symmetric manifolds, First Inter-

national Congress of Chinese Mathematicians (Beijing, 1998), AMS/IP Stud. Adv. Math., vol. 20, American Mathematical Society, Providence, Rhode Island, 2001, pp. 271–278.

A. Lichnerowicz, Les varié té s de Poisson et leurs algè bres de Lie associé es, J. Differential

Geometry 12 (1977), 253–300. G. Lion and M. Vergne, The Weil Representation, Maslov Index, and Theta Series, Birkhäuser,

Boston-Basel-Berlin, 1980. O. Loos, Symmetric Spaces, vol I, Benjamin, New York-Amsterdam, 1969. G. Lusztig, A class of irreducible representations of a Weyl group, Proc. Kon. Nederl. Akad., A

82 (1982), 323–335. G. Lusztig, Characters of Reductive Groups over a Finite Field, Annals of Mathematics Studies,

vol. 107, Princeton University Press, Princeton, New Jersey, 1984. G. Lusztig, Intersection cohomology complexes on a reductive group Invent. Math. 75 (1984),

205–272. G. Lusztig, Character sheaves I, Adv. in Math. 56 (1985), 193–237. G. Lusztig, Cuspidal local systems and graded Hecke algebras, Inst. Hautes Études Sci. Publ.

Math. 67 (1988), 145–202. G. Lusztig, Cells in affine Weyl groups, IV, J. Fac. Sci. Univ. Tokyo Sec. 1A Math. 36 (1989),

297–328. G. Lusztig, Affine Hecke algebras and their graded version, J. Amer. Math. Soc. 2 (1989), 599–635. G. Lusztig, Bases in equivariant K-theory, Represent. Theory 2 (1998), 298–369. G. Lusztig and N. Spaltenstein, Induced unipotent classes, J. London Math. Soc.(2) 19 (1979),

41–52. G. Lusztig and D. Vogan, Singularities of closures of K-orbits on flag manifolds, Invent. Math.

71 (1983), 365–379. I. Macdonald, Symmetric Functions and Hall Polynomials, Oxford Mathematical Monographs,

Clarendon Press, Oxford University Press, New York, 1995. G. Mackey, Mathematical Foundations of Quantum Mechanics, W. A. Benjamin, Inc., New York,

1963. G. Mackey, Theory of Unitary Group Representations, University of Chicago Press, Chicago,

1976. R. MacPherson, Chern classes for singular algebraic varieties Ann. of Math. 100 (1974), 423–432. A. I. Mal’cev, On semi-simple subgroups of Lie groups (Russian), Izvestia Akad. Nauk SSSR 8

(1944), 143–174. A. I. Mal’cev,, On semi-simple subgroups of Lie groups, Amer. Math. Soc. Transl., Ser. 1 11

(1950), 172–213. G. Margulis, Discrete Subgroups of Semisimple Lie Groups, Springer-Verlag, Berlin-Heidelberg-

New York, 1991. J. Marsden and A. Weinstein, Reduction of symplectic manifolds with symmetry, Rep. Mathe-

matical Phys. 5 (1974), 121–130. T. Matsuki, The orbits of affine symmetric spaces under the action of minimal parabolic sub-

groups, J. Math. Soc. Japan 31 (1979), 331–357. Y. Matsushima, On Betti numbers of compact locally symmetric Riemannian manifolds, Jour.

Diff. Geom. 1, 99–109. H. Matsumura, Commutative Ring Theory, translated by M. Reid, Cambridge University Press,

Cambridge, 1986. F. I. Mautner, Unitary representations of locally compact groups I, Ann. of Math. 51 (1950),

1–25.

F. I. Mautner, Unitary representations of locally compact groups II, Ann. of Math. 52 (1951),

528–556. W. McGovern, Rings of regular functions on nilpotent orbits and their covers, Invent. Math. 97

(1989), 209–217. W. McGovern, Unipotent representations and Dixmier algebras, Compositio math. 69 (1989),

241–276. W. McGovern, Completely Prime Maximal Ideals and Quantization, Memoirs of the AMS, Amer-

ican Mathematical Society, Providence, Rhode Island, 1994. W. McGovern, Rings of regular functions on nilpotent orbits II: model algebras and orbits, Comm.

Alg. 22 (1994), 765–772. J. Mickelsson, Step algebras of semi-simple subalgebras of Lie algebras, Rep. Mathematical Phys.

4 (1973), 307–318. D. Miličić, On C^{∗}-algebras with bounded trace, Glasnik Mat. 8 (28) (1973), 7–22. D. Miličić, On structure of dual of semisimple groups, Ph.D. dissertation, University of Zagreb

(1973). D. Miličić, The dual spaces of almost connected reductive groups, Glasnik Mat. 9 (29) (1974),

273–288. C. Moeglin, Idéaux primitifs complètement premiers de l’algèbre enveloppante 𝔤𝔩(n, ℂ), J./ Alg./

106 (1987), 287–366. C. Moeglin, Modèles de Whittaker et idéaux primitifs complètement premiers dans les algèbres

de lie semi-simples complexes II, Math. Scand. 63 (1988), 5-35. C. Moeglin, Orbites unipotentes et spectre discret non ramifié: le cas des groupes classiques

déployés, Compositio Math. 77 (1991), 1–54. C. Moeglin, M.-F. Vigneras, and J.-L. Waldspurger, Correspondances de Howe sur un Corps

p-adique, Lecture Notes in Mathematics, vol. 1291, Springer-Verlag, Berlin-Heidelberg-New York, 1987.

G. D. Mostow, Self-adjoint groups, Ann. of Math.(2) 62 (1955), 44–55. G. Muić, The unitary dual of p-adic G_{2}, Duke Math. J. 90 (1997), 465–493. B. J. Müller, Localization in non-commutative Noetherian rings, Canad. J. Math. 28 (1976),

600–610. D. Mumford, J. Fogarty, and F. Kirwan, Geometric Invariant Theory, Springer-Verlag, Berlin-

Heidelberg-New York, 1994. M. A. Naimark, On the description of all unitary representations of the complex classical groups

I, Mat. Sb. 35 (1954), 317–356. M. A. Naimark, On the description of all unitary representations of the complex classical groups

II, Mat. Sb. 37 (1955), 121–140. M. Nazarov, Representations of twisted Yangians associated with skew Young diagrams, preprint. E. Nelson, Analytic vectors, Ann. of Math. 70 (1959), 572–615. P. E. Newstead, Lectures on Introduction to Moduli Problems and orbit spaces, Springer-Verlag,

Berlin-Heidelberg-New York, 1978. T. Oshima and T. Matsuki, A description of discrete series for semisimple symmetric spaces,”,

Advanced Studies in Pure Mathematics, vol. 4: Group Representations and Systems of Differential Equations Kinokuniya, Tokyo, and North-Holland, Amsterdam-New York-Oxford,, 1984.

K. R. Parthasarathy, R. Ranga Rao, and V. S. Varadarajan, Representations of complex semi-

simple Lie groups and Lie algebras Ann. of Math. 85 (1967), 383–429. H. Ozeki and M. Wakimoto, On polarizations of certain homogeneous spaces, Hiroshima Math.

J. 2 (1972), 445–482. D. I. Panyushev, Rationality of singularities and the Gorenstein property of nilpotent orbits,

Functional Anal. Appl. 25 (1991), 225–226. R. Parthasarathy, Dirac operator and the discrete series, Ann. of Math. 96 (1972), 1–30. R. Parthasarathy, A generalization of the Enright-Varadarajan modules, Comp. Math. 36, 53–73. R. Parthasarathy, Criteria for the unitarizability of some highest weight modules Proc. Indian

Acad. Sci. 89 (1980), 1–24. F. Peter and H. Weyl, Die Vollständigkeit der primitiven Darstellungen einer geschlossenen kon-

tinuierlichen Gruppe, Math. Ann. 97 (1927), 737–755. Pontrjagin, Topological Groups, translated by E. Lehmer, Princeton University Press, Princeton,

New Jersey, 1939.

L. Pukanszky, The Plancherel formula for the universal covering group of SL(2, ℝ), Math. Ann.

156 (1964), 96–143. J. Rawnsley, W. Schmid, and J. Wolf, Singular unitary representations and indefinite harmonic

theory, J. Funct. Anal. 51 (1983), 1–114. M. Reeder, Zero weight spaces and the Springer correspondence, Indag. Math. (N.S.) 9 (1998),

431–441. R. Rentschler, Comportement de l’application de Dixmier par rapport à l’anti-automorphism

principal pour les algèbres de Lie résolubles, C. R. Acad. Sci. Paris 282 (1976), 555-557. R. Richardson, Conjugacy classes in parabolic subgroups of semi-simple algebraic groups, Bull.

London Math. Soc. 6 (1974), 21–24. M. Rosenlicht, Some basic theorems on algebraic groups, Amer. J. Math. 78 (1956), 401–443. L. Rothschild and J. Wolf, Representations of semisimple Lie groups associated to nilpotent coad-

joint orbits, Ann. Sci. École Norm. Sup. (4) 7 (1974), 155–173. W. Rudin, Real and Complex Analysis, McGraw-Hill, New York, 1966. S. Salamanca-Riba, On the unitary dual of some classical Lie groups, Compositio Math. 68

(1988), 251–303. S. Salamanca-Riba, On the unitary dual of real semisimple Lie groups snd the A_{𝔮}(λ) modules.

The regular case, Duke Math. J. 96 (1999), 521–546. S. Salamanca-Riba and D. Vogan, On the classification of unitary representations of reductive

Lie groups, Ann. of Math. 148 (1998), 1067–1133. M. Sato, M. Kashiwara, and T. Kawai, Hyperfunctions and pseudodifferential equations, Lecture

Notes in Mathematics, vol. 287, Springer-Verlag, Berlin-Heidelberg-New York,, 1973, pp. 265– 529.

W. Schmid, Homogeneous complex manifolds and representations of semisimple Lie groups,

Ph.D. dissertation, University of California at Berkeley, Berkeley, 1967. W. Schmid, Homogeneous complex manifolds and representations of semisimple Lie groups, Rep-

resentation Theory and Harmonic Analysis on Semisimple Lie Groups (P. Sally and D. Vogan, eds.), Mathematical Surveys and Monographs 31, American Mathematical Society, Providence, Rhode Island, 1989, pp. 223–286.

W. Schmid, Die Randwerte holomorpher Funktionen auf hermitesch symmetrischen Räumen,

Invent. Math. 9 (1969–1970), 61–80. W. Schmid, L_{2} cohomology and the discrete series, Ann. of Math. 103 (1976), 375–394. W. Schmid, Some properties of square-integrable representations of semisimple Lie groups, Ann.

of Math. 102 (1975), 535–564. W. Schmid, On the characters of the discrete series (the Hermitian symmetric case), Invent.

Math. 30 (1975), 47–144. W. Schmid, Two character identities for semisimple Lie groups, Non-commutative Harmonic

Analysis and Lie groups (J. Carmona and M. Vergne, eds.), Lecture Notes in Mathematics, vol. 587, Springer-Verlag, Berlin-Heidelberg-New York, 1977, pp. 196–225.

W. Schmid, Discrete series, These proceedings. W. Schmid and J. Wolf, A vanishing theorem for open orbits on complex flag manifolds, Proc.

Amer. Math. Soc. 92 (1984), 461–464. W. Schmid, Boundary value problems for group invariant differential equations, The mathematical

heritage of É lie Cartan, Astérisque, vol. 311–321, 1985. W. Schmid, Boundary value problems for group invariant differential equations, The mathematical

heritage of Élie Cartan, Astérisque, vol. 311–321, 1985. J. Schwartz, The determination of the admissible orbits in the real classical groups, Ph.D. disser-

tation, Massachusetts Institute of Technology, Cambridge, Massachusetts, 1987. J. Sekiguchi, Remarks on nilpotent orbits of a symmetric pair J. Math. Soc. Japan 39 (1987),

127–138. I. Shafarevich, Basic Algebraic Geometry; K. Hirsch, Springer-Verlag, Berlin, Heidelberg, New

York, 1977. I. E. Segal, An extension of Plancherel’s formula to separable unimodular groups, Ann. of Math.

52 (1950), 272–292.. I. E. Segal, Hypermaximality of certain operators on Lie groups, Proc. Amer. Math. Soc. 3 (1952),

13–15.

J. P. Serre, Représentations linéaires et espaces homogènes kählériens des groupes de Lie compacts, Exp. 100, Séminaire Bourbaki, 6ième année: 1953/54, Secrétariat mathématique, Paris, 1959.

J. P. Serre, Un théorème de dualité, Comment. Math. Helv. 29 (1955), 9–26. J. P. Serre, Algèbres de Lie Semi-Simples Complexes, W. A. Benjamin, New York, 1966. J. P. Serre, Local Fields, Springer-Verlag, New York, Heidelberg, Berlin, 1979. D. Shale, Linear symmetries of free boson fields, Trans. Amer. Math. Soc. 103 (1962), 149–167. I. M. Shchepochkina, The orbit method in restriction and induction problems for solvable Lie

groups, Selecta Math. 13 (1994), 305–355. D. Shelstad, Characters and inner forms of a quasi-split group over ℝ Compositio Math. 39

(1979), 11–45. D. Shelstad, L-indistinguishability for real groups Math. Ann. 259 (1982), 385–430. D. Shelstad, Orbital Integrals, Endoscopic Groups, and L-indistinguishability for Real Groups,

Publ. Math. de l’Université de Paris, 1983. A. Silberger, PGL_{2} over the p-adics: its representations, spherical functions, and Fourier analy-

sis, Lecture Notes in Mathematics, vol. 166, Springer-Verlag, Berlin-New York, 1970. W. Soergel, An irreducible not admissible Banach representation of SL(2, ℝ), Proc. Amer. Math. Soc.

104 (1988), 1322–1324. J.-M. Souriau, Structure des systèmes dynamiques, Maîtrises de mathématiques, Dunod, Paris,

1970; English translation by C. H. Cushman-de Vries in Structure of dynamical systems: A symplectic view of physics, Progress in Mathematics, vol. 149, Birkhuser Boston, Boston, 1997.

N. Spaltenstein, On the fixed point set of a unipotent element on the variety of Borel subgroups,

Topology 16 (1977), 203–204. B. Speh, Some results on principal series for GL(n, ℝ), Ph.D. dissertation, Massachusetts Insti-

tute of Technology, Cambridge, Massachusetts, June, (1977). B. Speh, Unitary representations of SL(n, ℝ) and the cohomology of congruence subgroups, Non-

commutative Harmonic Analysis and Lie groups (J. Carmona and M. Vergne, eds.), Lecture Notes in Mathematics, vol. 880, Springer-Verlag, Berlin-Heidelberg-New York,, 1981, pp. 483– 505.

B. Speh, The unitary dual of GL(3,ℝ) and GL(4,ℝ), Math. Ann. 258 (1981), 113–133. B. Speh and D. Vogan, Reducibility of generalized principal series representations, Acta Math.

145 (1980), 227–299. T. A. Springer, Reductive groups, Automorphic Forms, Representations, and L-functions, Pro-

ceedings of Symposia in Pure Mathematics, vol. 33, part 1, American Mathematical Society, Providence, Rhode Island,, 1979, pp. 3–28.

T. A. Springer, Microlocalisation algébrique, Séminaire d’algèbre Paul Dubreil et Marie-Paul Malliavin, 36ème année (Paris, 1983–1984), Lecture Notes in Mathematics, vol. 1146, Springer-Verlag, Berlin-New York, 1985, pp. 299–316.

T. A. Springer, Linear Algebraic Groups, Birkhäuser, Boston-Basel-Stuttgart, 1981. E. M. Stein, Analysis in matrix spaces and some new representations of SL(n,ℂ), Ann. of Math.

86 (1967), 461–490. R. Steinberg, Lectures on Chevalley Groups, Yale University, New Haven, Connecticut, 1968. D. Sternheimer, Phase-space representations,” Lectures in Applied Mathematics 21 (1985), Amer-

ican Mathematical Society, Providence, Rhode Island. M. Tadić, The topology of the dual space of a reductive group over a local field, Glasnik Mat. 18

(38) (1983), 259–279. M. Tadić, Classification of unitary representations in irreducible representations of general linear

group (non-Archimedean case), Ann. . Sci. École Norm. Sup.(4) 19 (1986), 335–382. M. Tadić, Topology of unitary dual of nonarchimedean GL(n), Duke Math. J. 55 (1987), 385–422. M. Tadić, Geometry of dual spaces of reductive groups (non-archimedean case), J. Analyse Math.

51 (1988), 139–181. J. Tate, Number theoretic background, Automorphic Forms, Representations, and L-functions,

Proceedings of Symposia in Pure Mathematics, vol. 33, part 2, American Mathematical Society, Providence, Rhode Island, 1979, pp. 3–26.

J. Tirao and J. Wolf, Homogeneous holomorphic vector bundles, Indiana Univ. Math. J. 20 (1970/1971), 15–31.

P. Torasso, Quantification géométrique, opérateurs d’entrelacement et représentations unitaires

de _{SL}_{3}(ℝ), Acta Math. 150 (1983), 153–242. P. Torasso, Méthode des orbites de Kirillov-Duflo et représentations minimales des groupes sim-

ples sur un corps local de caractéristique nulle, Duke Math. J. 90 (1997), 261–377. F. Treves, Topological vector spaces, distributions, and kernels, Academic Press, New York, 1967. M. Tsuchikawa, On the representations of SL(3,ℂ), III, Proc. Japan Acad. 44 (1968), 130–132. W. M. Tulczyjew, Poisson brackets and canonical manifolds, Bull. Acad. Polon. Sci. Sé r. Sci.

Math. Astronom. Phys. 22 (1974), 931–935. I. Vakhutinski, Unitary representations of GL(3,ℝ), Mat. Sb. 75 (1968), 303–320. G. van Dijk, Invariant eigendistributions on the tangent space of a rank one semisimple symmetric

space Math. Ann. 268 (1984), 405–416. V. S. Varadarajan, Harmonic Analysis on Real Reductive Groups, Lecture Notes in Mathematics,

vol. 576, Springer-Verlag, Berlin-Heidelberg-New York, 1977. V. S. Varadarajan, Lie groups, Lie algebras, and their representations, Graduate Texts in Math-

ematics, vol. 102, Springer-Verlag, New York-Berlin, 1984. M. Vergne, Instantons et correspondance de Kostant-Sekiguchi, C. R. Acad. Sci. Paris Sér. I Math.

320 (1995), 901–906. D. N. Verma, Structure of certain induced representations of complex semi–simple Lie algebras

dissertation, Yale University, (1966). D. Vogan, Gelfand-Kirillov dimension for Harish-Chandra modules, Invent. Math. 48 (1978),

75–98. D. Vogan, The algebraic structure of the representations of semisimple Lie groups I, Ann. of

Math. 109 (1979), 1–60. D. Vogan, Irreducible characters of semisimple Lie groups I, Duke Math. J. 46 (1979), 61–108. D. Vogan, Irreducible characters of semisimple Lie groups II. The Kazhdan-Lusztig conjectures,

Duke Math. J. 46 (1979), 805–859. D. Vogan, A generalized τ-invariant for the primitive spectrum of a semsimple Lie algebra, Math.

Ann. 242 (1979), 209–244. D. Vogan, Singular unitary representations, Non-commutative Harmonic Analysis and Lie groups

(J. Carmona and M. Vergne, eds.), Lecture Notes in Mathematics, vol. 880, Springer-Verlag, Berlin-Heidelberg-New York, 1981, pp. 506–535.

D. Vogan, Representations of Real Reductive Lie Groups, Birkhäuser, Boston-Basel-Stuttgart,

1981. D. Vogan, Irreducible characters of semisimple Lie groups III. Proof of the Kazhdan-Lusztig

conjectures in the integral case, Invent. Math. 71 (1983), 381–417. D. Vogan, Irreducible characters of semisimple Lie groups IV. Character-multiplicity duality,

Duke Math. J. 49 (1982), 943–1073. D. Vogan, Understanding the unitary dual, Lie Group Representations I (R. Herb, R. Lipsman

and J. Rosenberg, eds.), Lecture Notes in Mathematics, vol. 1024, Springer-Verlag, Berlin-Heidelberg-New York, 1983, pp. 264–286.

D. Vogan, Unitarizability of certain series of representations, Ann. of Math. 120 (1984), 141–187. D. Vogan, The unitary dual of GL(n) over an archimedean field, Invent. Math. 83 (1986), 449–

505. D. Vogan, The orbit method and primitive ideals for semisimple Lie algebras,”, Lie Algebras and

Related Topics (D. Britten, F. Lemire, and R. Moody, eds.), CMS Conference Proceedings, vol. 5, American Mathematical Society for CMS, Providence, Rhode Island, 1986.

D. Vogan, Representations of reductive Lie groups 245-266, Proceedings of the International Congress of Mathematicians 1986, volume I, American Mathematical Society, Providence, Rhode Island, 1987.

D. Vogan, Unitary Representations of Reductive Lie Groups, Annals of Mathematics Studies,

Princeton University Press, Princeton, New Jersey, 1987. D. Vogan, Irreducibility of discrete series representations for semisimple symmetric spaces, Rep-

resentations of Lie Groups, Kyoto, Hiroshima, 1986 (K. Okamoto and T. Oshima, eds.), Advanced Studies in Pure Mathematics, volume 14, Kinokuniya Company, Ltd., Tokyo, 1988, pp. 191–221.

D. Vogan, Noncommutative algebras and unitary representations, The Mathematical Heritage of Hermann Weyl (R. O. Wells, Jr., ed.), Proceedings of Symposia in Pure Mathematics, vol. 48, American Mathematical Society, Providence, Rhode Island, 1988.

D. Vogan, Dixmier algebras, sheets, and representation theory, Operator Algebras, Unitary Representations, Enveloping Algebras, and Invariant Theory (A. Connes, M. Duflo, A. Joseph, and R. Rentschler, eds.), Birkhäuser, Boston, MA, 1990, pp. 333-395.

D. Vogan, Associated varieties and unipotent representations, Harmonic Analysis on Reductive

Groups (W. Barker and P. Sally, eds.), Birkhäuser, Boston-Basel-Berlin, 1991, pp. 315–388. D. Vogan, Unitary representations of reductive Lie groups and the orbit method, New Develop-

ments in Lie Theory and Their Applications (J. Tirao and N. Wallach, eds.), Birkhäuser, Boston-Basel-Berlin, 1992, pp. 87–114.

D. Vogan, The local Langlands conjecture, Representation Theory of Groups and Algebras (J. Adams et al., eds.), Contemporary Mathematics 145, American Mathematical Society, Providence, Rhode Island, 1993, pp. 305–379.

D. Vogan, The orbit method and unitary representations of reductive Lie groups, Algebraic and Analytic Methods in Representation Theory (Sønderborg, 1994), Perspectives in Mathematics, vol. 17, Academic Press, San Diego, 1997, pp. 243–339.

D. Vogan, Unipotent representations and cohomological induction, The Penrose Transform and Analytic Cohomology in Representation Theory (M. Eastwood, J. Wolf, and R. Zierau, eds.), Contemporary Mathematics 154, American Mathematical Society, Providence, Rhode Island, 1993, pp. 47–70.

D. Vogan, The unitary dual of G_{2}, Invent. Math. 116 (1994), 677–791. D. Vogan, Cohomology and group representations, Representation Theory and Automorphic Forms

(Instructional Conference, International Centre for Mathematical Sciences, Edinburgh, March, 1996) (T. Bailey and A. Knapp, eds.), Proceedings of Symposia in Pure Mathematics, vol. 61, American Mathematical Society, Providence, RI, 1997, pp. 219–243.

D. Vogan, A Langlands classification for unitary representations, Analysis on Homogeneous spaces and Representations of Lie groups, Advanced Studies in Pure Mathematics, volume 26, Kinokuniya Company, Ltd., Tokyo, 1998.

D. Vogan, The method of coadjoint orbits for real reductive groups, Representation Theory of Lie Groups, IAS/Park City Mathematics Series, vol. 8, American Mathematical Society, Providence, RI, 1999.

D. Vogan, Unitary representations of reductive Lie groups, Mathematics Towards the Third Mil-

lenium, Rend. Mat. Acc. Lincei s. 9, fasc. spec., Rome, 2000, pp. 147–167. D. Vogan and N. Wallach, Intertwining operators for real reductive groups, Adv. in Math. 82

(1990), 203–243. D. Vogan and G. Zuckerman, Unitary representations with non-zero cohomology, Compositio

Math. 53 (1984), 51–90. J. von Neumann, Die eindeutigkeit der Schrödingerschen Operatoren, Math. Ann. 104 (1931),

570–578. T. Vust, Opèration de groupes réductifs dans un type de cônes presques homogènes Bull. Soc.

Math. France 102 (1974), 317–333. N. Wallach, Harmonic Analysis on Homogeneous Spaces, Marcel Dekker, New York, 1973. N. Wallach, On the unitarizability of derived functor modules, Invent. Math. 78 (1984), 131–141. N. Wallach, Real Reductive Groups I, Academic Press, San Diego, 1988. N. Wallach, Real Reductive Groups II, Academic Press, San Diego, 1992. F. Warner, Foundations of Differentiable Manifolds and Lie Groups, Scott, Foresman and Com-

pany, Glenview, IL, 1971. G. Warner, Harmonic Analysis on Semisimple Lie Groups, vols. I and II, Springer-Verlag, Berlin,

Heidelberg, New York, 1972. A. Weil, L’intégration dans les Groupes Topologiques et ses Applications Actual. Sci. Ind., vol. 869,

Hermann et Cie., Paris, 1940. A. Weil, Sur certaines groupes d’opérateurs unitaires Acta Math. 111 (1964), 143–211. A. Weil, Basic Number Theory, Springer-Verlag, New York, Heidelberg, Berlin, 1974. A. Weinstein, The local structure of Poisson manifolds, J. Differential Geom. 18 (1983), 523–557. H. Weyl, Theorie der Darstellung kontinuierlicher Gruppen. I., Math. Z. 23 (1925), 271–309; II

24 (1925), 328–376; III (1925), 377–395. H. Weyl, The Theory of Groups and Quantum Mechanics, Dover, 1931. H. Weyl, The Classical Groups. Their Invariants and Representations, Princeton University Press,

Princeton, New Jersey, 1939.

E. Wigner, On unitary representations of the inhomogeneous Lorentz group, Ann. of Math. 40

(1939), 149–204. J. Wolf, The action of a real semisimple group on a complex flag manifold. I. Orbit structure and

holomorphic arc components, Bull. Amer. Math. Soc. 75 (1969), 1121–1237. H. Wong, Dolbeault cohomologies and Zuckerman modules associated with finite rank representa-

tions, Ph.D. dissertation, Harvard University, Cambridge, Massachusetts, 1992. H. Wong, Dolbeault cohomological realization of Zuckerman modules associated with finite rank

representations, J. Funct. Anal. 129 (1995), 428–454. H. Wong, Cohomological induction in various categories and the maximal globalization conjecture,

Duke Math. J. 96 (1999), 1–27. A. Zelevinsky, Induced representations of reductive 𝔭-adic groups II. On irreducible representa-

tions of GL(n) Ann. Sci. École Norm. Sup.(4) 13 (1980), 165–210. A. Zelevinsky, A p-adic analogue of the Kazhdan-Lusztig conjecture Funct. Anal. Appl. 15 (1981),

9–21. A. Zelevinsky, Two remarks on graded nilpotent classes Russian Math. Surveys 40 (1985), 249–

250. D. Zhelobenko, Harmonic Analysis on Complex Semisimple Lie Groups, Mir, Moscow, 1974. D. Želobenko, Compact Lie Groups and Their Representations, Translations of Mathematical

Monographs 40, American Mathematical Society, Providence, Rhode Island, 1973. D. P. Zhelobenko, Minimal K-types and classification of irreducible representations of reductive

Lie groups, Funktsional. Anal. i Prilozhen. 18 (1984), 79–80 (Russian); English transl. in Functional Anal. Appl. 18 (1984), 333–335.

R. Zierau, Unitarity of certain Dolbeault cohomology representations, The Penrose Transform and Analytic Cohomology in Representation Theory (M. Eastwood, J. Wolf, and R. Zierau, eds.), Contemporary Mathematics, vol. 154, American Mathematical Society, Providence, Rhode Island, 1993, pp. 239–259.

G. Zuckerman, Tensor products of finite and infinite dimensional representations of semisimple Lie groups, Ann. of Math. 106 (1977), 295–308.